OpenStudy (anonymous):

find the 50th term of sequence: -2, -9, -16 ... help

OpenStudy (anonymous):

what is the first term?

OpenStudy (anonymous):

do you know the first term?

OpenStudy (anonymous):

it is an A.P. c.d d=-9-(-2)=-9+2=-7 first term a=-2 nth term\[tn=a+(n-1)d\] find t50,here n=50